The perimeter of a rectangle is 38m. The length is 10m more than twice the width. Find the dimensions.
Answer by chriswen(106)

Let x m be the width.
Let 2x+10 m be the length.
...
P=2(l+w)
38=2(2x+10+x)
38=2(3x+10)
38=6x+20
6x=38-20
6x=18
x=18/6
x=3
2x+10=16
Therefore, the rectangle is 3m by 16m.